OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi divided by five.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you want a positive one, add \[\frac{\pi}{5}+2\pi\]

OpenStudy (misty1212):

you want a negative one, subtract !

OpenStudy (anonymous):

so would it be pi/5+360 and pi/5-360?